Mature Conglomerate Corporation (MCC) just paid a dividend of $1.35 per share, and that dividend is expected to grow at a constant rate of 6.00% per year in the future. The company's beta is 1.55, the required return on the market is 12.50%, and the risk-free rate is 4.00%. What is the intrinsic value for MCC’s stock? Enter your answer rounded to two decimal places. Do not enter $ or comma in the answer box. For example, if your answer is $12,300.456 then enter as 12300.46 in the answer box.
Answer #1

required return= risk-free rate +Beta*(market rate- risk-free rate )

=4+1.55*(12.5-4)

=17.175%

Intrinsic value=D1/(Required return-Growth rate)

=(1.35*1.06)/(0.17175-0.06)

which is equal to

=12.81(Approx).